# Quantum Error Correction resilient against Atom Loss

### v1

@@ 1 Introduction
Neutral atom quantum processors also present noise sources that are specific to the platform, including leakage out of the computational subspace through, e.g. spontaneous decay from a highly-excited electronic Rydberg state to a lower-energy state, population remaining in the Rydberg state after a given quantum gate or black-body radiation [22]. To address leakage errors, hardware-specific approaches have been suggested [21] as well as circuit-based techniques [25]. For example, in alkaline-earth-like atoms such as 171Yb, encoding the computational states into metastable states causes a large fraction of gate errors to leak out of the qubit subspace. These leakage errors can be detected and converted to erasure errors, easier to decode than standard Pauli errors  [26, 22, 27].
@@
One particularly important source of errors that is specific to the neutral atom platform is the loss of atoms during computation. Atom loss can occur due to several mechanisms, including heating that causes atoms to escape the optical traps, collisions with background gas particles, or anti-trapping during an excitation to a Rydberg state [28, 25, 10, 29, 30]. We note that Rydberg leakage errors can also be converted into atom loss errors using the repulsive force of the trap [29]. A promising approach to identify and correct atom losses is provided by so-called loss detection units (LDU) [2, 4, 31, 25, 32, 29], which are small circuits attached to each data qubit, entangling the latter with an additional atom. The performances of QEC protocols supplemented with theses LDUs in the presence of atom loss have not been evaluated to date.
@@
In this work, we investigate QEC protocols supplemented with two types of LDUs: a standard LDU [2, 4], which operates by entangling the LDU’s atom with the presence of the data qubit, and a recently introduced teleportation-based LDU scheme [29] where the state of the original data atom is transferred to a new, fresh atom. We focus on the rotated surface code [33, 34, 35, 12], which has emerged as a leading choice for fault-tolerant quantum computation. We investigate its tolerance to atom loss errors and depolarizing noise, a common approach for modelling noisy gates, by performing circuit-level simulations. We derive the scaling laws for logical error probability in the asymptotic regime of low physical error probability. We demonstrate the existence of an error threshold that depends linearly on the probabilities of both atom loss and depolarizing noise. In particular, for zero depolarizing noise, we find an atom loss threshold around $2.6\%$ for both LDU schemes. Finally, we conduct a comparative analysis of both LDU protocols, finding that the teleportation-based LDU outperforms the standard LDU in terms of achievable logical error probabilities, since it requires fewer two-qubit gates. It also avoids the need for active feedback, however at the cost of increased atom resources.
@@
For our analysis, we extend the minimal weight perfect matching (MWPM) decoder [36, 37, 38], a widely used decoding algorithm, to handle losses. Specifically, we incorporate the available information about the loss locations, namely the rounds and positions of lost qubits, into the algorithm by modifying the graph’s weights on which MWPM is executed. In the absence of depolarizing noise, this allows us to correct up $d-1$ errors for a surface code of distance $d$, analogous to an erasure channel. The decoding algorithm leads to a logical error probability reduction of almost three orders of magnitude compared to a naive decoding scheme.
@@
This article is organized as follows: Section 2 outlines the error models used in our study, including both atom loss and depolarizing noise. Section 3 introduces the two distinct LDU protocols — the standard LDU and the teleportation-based LDU — detailing their design and operations. In Section 4, we present the principle of the new decoder adapted for losses. Using this decoding scheme, we analyze in Section 5 the logical error probabilities for both protocols under varying noise conditions. We further discuss the results and compare the performance of the protocols. To emphasize the performance of our decoding scheme, we compare it to a naive decoder in Section 6. Finally, in Section 7, we detail the numerical method used to implement the surface code supplemented with these LDUs.
@@ 2 Error models
In this section, we detail the circuit-level error models used to evaluate the performance of QEC. In Section 2.1, we first describe the atom loss model which accounts for losses during gate operations due to Rydberg excitation and other mechanisms such as collisions and heating effects. We then present the depolarizing noise model, which characterizes imperfections arising from pulse applications and environmental interactions in Section 2.2. These models are combined to provide a comprehensive framework for simulating the impact of errors on neutral atom systems.
@@ 2.1 Atom loss
In our work, we assign to each two-qubit CZ gate a certain probability $p_{l}$ of losing an atom and assume that if both atoms are lost during the gate, the two losses can be described as two independent events (leading to a joint loss probability of $p_{l}^{2}$). We note that this approximation overestimates the loss probability compared to experiments, where a perfect blockade mechanism would prevent both atoms from being simultaneously excited to the Rydberg state, and thus to be simultaneously lost during the same CZ gate. Furthermore, when an atom is lost during a CZ gate, it is assumed that no entangling operation is carried out on the remaining atom. Any subsequent operation involving the lost atom (both one-qubit and two-qubit operations) is erased until a new atom has been reloaded (see Fig. 1).
@@ 3.2 Teleportation-based LDU
The teleportation-based LDU offers several advantages over the standard protocol. First, it requires only one CZ gate compared to two in the standard protocol. Second, it eliminates the need for active feedback gate during computation, whereas the standard protocol necessitates different actions depending on the ancilla’s state. Finally, replacing all data qubits every round helps to cool down the system, thereby reducing losses due to heating. Although this cooling advantage is not considered in our simulation, it remains a significant benefit [29]. The main drawback of this scheme is its higher atom consumption.
@@
Our decoder is based on the MWPM algorithm [36, 37, 38] to decode errors resulting from depolarizing noise and atom loss. Error mechanisms generate a detector error model (DEM) consisting of a set of lists of flipped stabilizers along with their probabilities of occurrence. Using Pymatching [38], a Python library that implements the MWPM algorithm, a graph is generated where nodes are the stabilizers and the edges represent error mechanisms pairing the corresponding flipped stabilizers. Each edge is weighted by the log-likelihood ratio $\text{ln}(\frac{1-p}{p})$ where $p$ is the probability of the error [34, 38]. This graph is extended in the time direction to account for multiple rounds of error correction, enabling us to decode errors stemming from stabilizer measurements [35].
@@
To achieve this, we augment the DEM generated by the depolarizing noise channel with new error mechanisms arising from atom loss. The absence of an atom at the end of a round $r$ results in a list of potential loss locations. Each potential loss location generates its own DEM. The probabilities of these DEMs are then weighted by the likelihood that the loss occurred at that specific location, conditioned that a loss was detected in the round $r$. These weighted DEMs are subsequently added to the final DEM, allowing the decoder to account for both depolarizing noise and atom loss in the error correction process.
@@ 5.4 Comparison of both LDU protocols
We plot the gain $G_{\text{LDU}}^{\text{decoder}}$ for both LDU protocols in Fig. 9 in the case of a surface code of distance $d=11$. One important outcome of this study is that the “loss-aware” decoder significantly outperforms the naive decoder for both LDU protocols, particularly below the threshold and when $p_{l}\lesssim p_{d}$. The maximum gain reaches comparable values of up to $2.8$ for the teleportation LDU protocol and $2.7$ for the standard LDU protocol, for sufficiently small $p_{d}$.
@@ 7.1.1 Loss probabilities for stabilizer operations
First, we compute the probability to lose an atom during the stabilizer operations. The loss probability of each individual CZ gate is denoted as $p_{l}$. For each atom, the probability to be lost at the $i$-th CZ gate in a round is given by Eq. (1).
@@ 7.1.3 Simulation of the standard LDU
Figure 11: Operations seen by a bulk data atom at each cycle when the surface code is equipped with the standard LDU. The original circuit with the standard LDU (top left circuit) is never simulated. In the absence of atom loss (top right circuit), the action of the standard LDU is trivial. To simulate atom loss, gates are erased in the circuit. Crosses indicate the potential loss locations in the original circuit. The arrows show the effective circuits being simulated on Stim along with their probabilities $p_{l,i}$ (see Eq. (1) for $0<i\leq 4$ and Eq. (10) and Eq. (12) for $i=n^{\text{stab.}}_{\text{CZ}}+1,n^{\text{stab.}}_{\text{CZ}}+2$). $n_{\text{ancilla lost}}$ is the number of time the LDU’s ancilla is lost (its potential loss locations are indicated by green crosses) during the standard LDU. This loss can be detected and another LDU is reapplied in that case. Red boxes represents the depolarizing noise channel associated with CZ gates. The resulting 1-qubit effective noise channel stemming from the noisy standard LDU has an error probability $p_{d_{1}}^{\text{stand. LDU}}$ (see Eqs. (19) (20)) , which we denote simply as $p_{d_{1}}$ in the figure to avoid visual clutter. Noise in the standard LDU also induces a flip with probability $p_{\text{flip}}$ (see Eq. (23)) in the measurement outcome causing the loss data atom not being reloaded (red crosses) or reloading an new atom while still being present (red |$\ket{0}$ on the top right circuit).
@@ 7.2.1 Effective 1-qubit depolarizing channel
$f_{d}^{\text{stand. LDU}}$ $=\sum_{r=0}^{\infty}(1-(1-p_{l})^{2})^{r}\left(1-\frac{16p_{d}}{15}\right)^{2r+2}$ (18)
$=\frac{\left(1-\frac{16p_{d}}{15}\right)^{2}}{1-(1-(1-p_{l})^{2})(1-\frac{16p_{d}}{15})^{2}}\quad.$ (19)
@@
$p_{d_{1}}^{\text{stand. LDU}}=\frac{3}{4}\frac{1-(2-(1-p_{l})^{2})(1-\frac{16p_{d}}{15})^{2}}{1-(1-(1-p_{l})^{2})(1-\frac{16p_{d}}{15})^{2}}\quad.$ (20)
@@ 7.2.2 Measurement flips
$p_{\text{flip}}$ $=\frac{1-f_{d}^{\text{stand. LDU}}}{2}$ (22)
$=\frac{1-(1-\frac{16p_{d}}{15})^{2}}{2}\quad.$ (23)

### v2

@@ 1 Introduction
Neutral atom quantum processors also present noise sources that are specific to the platform, including leakage out of the computational subspace through, e.g. spontaneous decay from a highly-excited electronic Rydberg state to a lower-energy state, population remaining in the Rydberg state after a given quantum gate or black-body radiation [22]. To address leakage errors, hardware-specific approaches have been suggested [21, 25] as well as circuit-based techniques [26]. For example, in alkaline-earth-like atoms such as 171Yb, encoding the computational states into metastable states causes a large fraction of gate errors to leak out of the qubit subspace. These leakage errors can be detected and converted to erasure errors, easier to decode than standard Pauli errors  [27, 22, 28, 29, 25]. Erasure conversion mechanism is not limited to neutral atom architectures and has also been studied in the context of superconducting circuits [30, 31, 32, 33].
@@
One particularly important source of errors that is specific to the neutral atom platform is the loss of atoms during computation. Atom loss can occur due to several mechanisms, including heating that causes atoms to escape the optical traps, collisions with background gas particles, or anti-trapping during an excitation to a Rydberg state [34, 26, 10, 35, 36]. We note that Rydberg leakage errors can also be converted into atom loss errors using the repulsive force of the trap [35]. A promising approach to identify and correct atom losses is provided by so-called loss detection units (LDU) [2, 4, 37, 30, 38, 26, 39, 35], which are small circuits attached to each data qubit, entangling the latter with an additional atom. The performances of QEC protocols supplemented with theses LDUs in the presence of atom loss have not been evaluated to date. We note that a recent independent work [29] examines atom loss in the context of measurement-based quantum computation. Employing a 3D RHG cluster state decoder, they reaches similar results, consistent with our work.
@@
In this work, we investigate QEC protocols supplemented with two types of LDUs: a standard LDU [2, 4], which operates by entangling the LDU’s atom with the presence of the data qubit, and a recently introduced teleportation-based LDU scheme [35] where the state of the original data atom is transferred to a new, fresh atom. We focus on the rotated surface code [40, 41, 42, 12], which has emerged as a leading choice for fault-tolerant quantum computation. We investigate its tolerance to atom loss errors and depolarizing noise, a common approach for modelling noisy gates, by performing circuit-level simulations. We derive the scaling laws for logical error probability in the asymptotic regime of low physical error probability. We demonstrate the existence of an error threshold that depends linearly on the probabilities of both atom loss and depolarizing noise. In particular, for zero depolarizing noise, we find an atom loss threshold around $2.6\%$ for both LDU schemes. Finally, we conduct a comparative analysis of both LDU protocols, finding that the teleportation-based LDU outperforms the standard LDU in terms of achievable logical error probabilities, since it requires fewer two-qubit gates. It also avoids the need for active feedback. Both schemes require approximately the same number of atoms, with a slight advantage for the teleportation-based LDU.
@@
For our analysis, we extend the minimal weight perfect matching (MWPM) decoder [43, 44, 45, 46], a widely used decoding algorithm, to handle losses. Specifically, we incorporate the available information about the loss locations, namely the rounds and positions of lost qubits, into the algorithm by modifying the graph’s weights on which MWPM is executed. In the absence of depolarizing noise, this allows us to correct up $d-1$ errors for a surface code of distance $d$, analogous to an erasure channel. The decoding algorithm leads to a logical error probability reduction of almost three orders of magnitude compared to a naive decoding scheme. We note that an independent study [47] that adopts the same decoding strategy was released shortly after our article. Additionally, another recent work [48] introduced an adaptive decoder called the local cluster decoder, which builds upon the Union-Find decoder. It enables efficient decoding of Pauli noise in the presence of leakage in the surface code.
@@
This article is organized as follows: Section 2 outlines the error models used in our study, including both atom loss and depolarizing noise. Section 3 introduces the two distinct LDU protocols — the standard LDU and the teleportation-based LDU — detailing their design and operations. In Section 4, we present the principle of the new decoder adapted for losses. Using this decoding scheme, we analyze in Section 5 the logical error probabilities for both protocols under varying noise conditions. We further discuss the results and compare the performance of the protocols. To emphasize the performance of our decoding scheme, we compare it to a naive decoder in Section 6. In Section 7, we detail the numerical method used to implement the surface code supplemented with these LDUs. Finally, in Section 8, we investigate a refined loss model that impact the neighbouring atom.
@@ 2 Error models
In this section, we detail the circuit-level error models used to evaluate the performance of QEC. In Sec. 2.1, we first describe the atom loss model which accounts for losses during gate operations due to Rydberg excitation and other mechanisms such as collisions and heating effects. We then present the depolarizing noise model, which characterizes imperfections arising from pulse applications and environmental interactions in Sec. 2.2. These models are combined to provide a comprehensive framework for simulating the impact of errors on neutral atom systems.
@@ 2.1 Atom loss
In our work, we assign to each two-qubit CZ gate a certain probability $p_{l}$ of losing an atom and assume that if both atoms are lost during the gate, the two losses can be described as two independent events (leading to a joint loss probability of $p_{l}^{2}$).Furthermore, when an atom is lost during a CZ gate, it is assumed that no entangling operation is carried out on the remaining atom. Any subsequent operation involving the lost atom (both one-qubit and two-qubit operations) is erased until a new atom has been reloaded (see Fig. 1).
We note that in principle the remaining atom may be impacted by the loss of the other atom during the CZ gate, possibly leading to a modified noise channel. In Sec. 8, we explore a refined loss model where the neighboring atom experiences a maximally biased Z error.
@@ 3.2 Teleportation-based LDU
3.3 Qualitative comparison of both protocols and atom consumption
The teleportation-based LDU offers several advantages over the standard protocol. First, it requires only one CZ gate compared to two in the standard protocol. Second, it eliminates the need for active feedback gate during computation, whereas the standard protocol necessitates different actions depending on the ancilla’s state. Finally, replacing all data qubits every round helps to cool down the system, thereby reducing losses due to heating. Although this cooling advantage is not considered in our simulation, it remains a significant benefit [35].
Assuming the ancilla LDU for the standard LDU protocol and the older data atom for the teleportation LDU can be reused in subsequent rounds (provided they are not lost), the atom consumption of the teleportation LDU exhibits a slight advantage over the standard LDU. Both protocols require trapping $d^{2}$ additional atoms (one for each data qubit) compared to the standard surface code (i.e. without LDU). During each error correction cycle, an average of $p_{l,\text{cycle}}\times n_{\text{atoms}}$ atoms are lost, where $p_{l,\text{cycle}}=1-(1-p_{l})^{n_{\text{CZ}}}\simeq p_{l}\times n_{\text{CZ}}$ represents the probability of losing an atom within a single round. The number of CZ gates experienced by an atom depends on its position in the surface code. For simplicity and to provide an estimate of atom consumption, we focus on atoms in the bulk, which are far more numerous, and neglect those located on the edges.
In both protocols, bulk ancilla atoms used for stabilizer measurements experience four CZ gates per cycle, while bulk data atoms experience six CZ gates per cycle. For the teleportation LDU, fresh atoms are not replaced in the same round as the older ones if they are lost. Thus, the atom loss only accounts for older data atoms, where we take into accounts the LDU operations acting on the data qubit during both the previous and current rounds. Consequently, the average atom consumption per cycle for the teleportation LDU scales as:
$n_{\text{tel. LDU}}\sim 10\times p_{l}\times d^{2}$ (2)
Additionally, the standard LDU requires considering the replacement of the ancilla LDU. They experience approximately two CZ gates per round (neglecting that we reapply a new LDU whenever the ancilla LDU is lost). As a result, the average number of atoms that need to be reloaded per cycle for the standard LDU scales as:
$n_{\text{stand. LDU}}\sim 12\times p_{l}\times d^{2}.$ (3)
@@ 3.3 Qualitative comparison of both protocols and atom consumption
Our decoder is based on the MWPM algorithm [43, 44, 45, 46] to decode errors resulting from depolarizing noise and atom loss. Error mechanisms generate a detector error model (DEM) consisting of a set of lists of flipped stabilizers along with their probabilities of occurrence. Using Pymatching [45, 46], a Python library that implements the MWPM algorithm, a graph is generated where nodes are the stabilizers and the edges represent error mechanisms pairing the corresponding flipped stabilizers. Each edge is weighted by the log-likelihood ratio $\text{ln}(\frac{1-p}{p})$ where $p$ is the probability of the error [41, 45, 46]. This graph is extended in the time direction to account for multiple rounds of error correction, enabling us to decode errors stemming from stabilizer measurements [42].
@@
To achieve this, we augment the DEM generated by the depolarizing noise channel with new error mechanisms arising from atom loss. The absence of an atom at the end of a round $r$ results in a list of potential loss locations. Each potential loss location generates its own DEM. The probabilities of these DEMs are then weighted by the likelihood that the loss occurred at that specific location, conditioned that a loss was detected in the round $r$. These weighted DEMs are subsequently added to the final DEM, allowing the decoder to account for both depolarizing noise and atom loss in the error correction process. Our decoder is similar to the one developped in [30] for the case of a SWAP-based LDU and a different loss model. Our approach offers the advantage of being easily automated using the DEM method implemented in Stim.
@@
Finally, the edge weights of the final DEM are defined as the sum of the independent loss events. This approach makes our decoder quite efficient, as we can precompute and store the DEM for each potential loss location before running the simulation. During the decoding process, we simply retrieve these precomputed DEMs whenever a loss at the corresponding location is detected.
@@ 5.4 Comparison of both LDU protocols
We plot the gain $G_{\text{LDU}}^{\text{decoder}}$ for both LDU protocols in Fig. 9 in the case of a surface code of distance $d=11$. One important outcome of this study is that the “loss-aware” decoder significantly outperforms the naive decoder for both LDU protocols, particularly below the threshold and when $p_{d}\lesssim p_{l}$. The maximum gain reaches comparable values of up to $2.8$ for the teleportation LDU protocol and $2.7$ for the standard LDU protocol, for sufficiently small $p_{d}$.
@@ 7.1.1 Loss probabilities for stabilizer operations
First, we compute the probability to lose an atom during the stabilizer operations. In our simulation, we first apply the Z stabilizers in parallel followed by the X stabilizers. The loss probability of each individual CZ gate is denoted as $p_{l}$. For each atom, the probability to be lost at the $i$-th CZ gate in a round is given by Eq. (1).
@@ 7.1.3 Simulation of the standard LDU
Figure 11: Operations seen by a bulk data atom at each cycle when the surface code is equipped with the standard LDU. The original circuit with the standard LDU (top left circuit) is never simulated. In the absence of atom loss (top right circuit), the action of the standard LDU is trivial. To simulate atom loss, gates are erased in the circuit. Crosses indicate the potential loss locations in the original circuit. The arrows show the effective circuits being simulated on Stim along with their probabilities $p_{l,i}$ (see Eq. (1) for $0<i\leq 4$ and Eq. (12) and Eq. (14) for $i=n^{\text{stab.}}_{\text{CZ}}+1,n^{\text{stab.}}_{\text{CZ}}+2$). $n_{\text{ancilla lost}}$ is the number of time the LDU’s ancilla is lost (its potential loss locations are indicated by green crosses) during the standard LDU. This loss can be detected and another LDU is reapplied in that case. Red boxes represents the depolarizing noise channel associated with CZ gates. The resulting 1-qubit effective noise channel stemming from the noisy standard LDU has an error probability $p_{d_{1}}^{\text{stand. LDU}}$ (see Eqs. (21), (22)) , which we denote simply as $p_{d_{1}}$ in the figure to avoid visual clutter. Noise in the standard LDU also induces a flip with probability $p_{\text{flip}}$ (see Eq. (24)) in the measurement outcome causing the loss data atom not being reloaded (red crosses) or reloading an new atom while still being present (red |$\ket{0}$ on the top right circuit).
@@ 7.2.1 Effective 1-qubit depolarizing channel
$f_{d}^{\text{stand. LDU}}$ $=\sum_{r=0}^{\infty}(1-(1-p_{l})^{2})^{r}\left(1-\frac{16p_{d}}{15}\right)^{2r+2}$
×$\times(1-p_{l})^{2}$ (20)
$=\frac{\left(1-\frac{16p_{d}}{15}\right)^{2}(1-p_{l})^{2}}{1-(1-(1-p_{l})^{2})(1-\frac{16p_{d}}{15})^{2}}\quad.$ (21)
@@
$p_{d_{1}}^{\text{stand. LDU}}=\frac{3}{4}\frac{1-(1-\frac{16p_{d}}{15})^{2}}{1-(1-(1-p_{l})^{2})(1-\frac{16p_{d}}{15})^{2}}\quad.$ (22)
@@ 7.2.2 Measurement flips
$p_{\text{flip}}$ $=\frac{1-(1-\frac{16p_{d}}{15})^{2}}{2}\quad.$ (24)
8 Loss with maximally biased Z error
In this final section, we modify the loss model and evaluate the resulting performance of the surface code. Experimentally, atomic loss caused by the Rydberg anti-trapping effect (or leakage outside the computational state) can occur at any time during the pulse application of the CZ gate. When this happens, the remaining atom is projected into the |$\ket{1}$ state (due to the Rydberg blockade). The remainder of the pulse subsequently re-excites the remaining atom, leaving it in the Rydberg state, which then decays into a state outside the computational subspace with high probability . We assume that such states can be immediately detected (e.g., via fluorescence ), enabling the repumping of the leaked atom back into the computational |$\ket{1}$ state. Using RC, this noise channel can be effectively transformed into a maximally (i.e. error probability of 0.5) biased Z-error channel  . Therefore, our revised loss model is the following: whenever an atom is lost during a given CZ gate, the neighboring remaining atom will experience a Z error of probability $0.5$.
In this section, we focus on the teleportation LDU as it shows the best performance for the simple loss model. Moreover, we anticipate the standard LDU to perform even worse than before compared to the teleportation LDU under this new loss model. This is because the conditional reapplication of the standard LDU, based on the loss detection of its ancilla increases the probability of making a Z error on the adjacent data qubit.
Figure 12: Performance analysis of the teleportation LDU scheme under the revised loss model. a) The logical error probability, normalized by the number of rounds for code distances $d=3,5,7,9,11$ as a function of the atom loss probability $p_{l}$ and at vanishing depolarizing noise $p_{d}=0$ (solid lines). The red vertical dashed line indicates the error threshold at $2.1\%$. Other color dashed lines show fitted curves using Eq. (LABEL:). For comparison, dotted lines represent results from for the simpler loss model studied in the rest of the article (see Figure 4b). b) Two dimensional color map of the logical error probability, normalized by the number of rounds for distance $d=11$ as a function of the loss probability $p_{l}$ and the depolarizing error probability $p_{d}$. The red solid line marks the error threshold while black solid lines show curves of constant logical error probability. The horizontal and vertical red dashed lines indicate the threshold at vanishing depolarizing noise and loss probability, respectively. Statistics were gathered using $10^{5}$ shots, except for the region below the $10^{-5}$ logical error probability contour in b) where $10^{6}$ shots were employed.
Fig. 12 presents the results obtained using the revised loss model. At vanishing depolarizing noise $p_{d}=0$ (see Fig. 12 a)), we observe a loss threshold of $2.1\%$, which is lower than the $2.6\%$ threshold found with the simpler loss model. For loss rates below $1\%$, the logical error probability per round is approximately one order of magnitude higher than in the simplified loss model (shown by dotted lines in Fig. 12a)). Since qubit loss affects neighboring qubits in this model, we expect the power law scaling exponent to be $(d+1)/2$, compared to an exponent of $d$ in the simpler loss model. We fit our numerical results using the following function:
$\sum_{k=\frac{d+1}{2}}^{d}a_{k}(d)p_{l}^{k}$ (25)
where $a_{k}(d)$ are fitting parameters that represents the number of weight-k physical errors that lead to a logical error for a code distance $d$. For the experimentally relevant parameter range ($0.001\leq p_{l}\leq 0.01$), our fitting analysis reveals an unexpected result: the dominant terms in Eq. (25) correspond to exponent close to $d$ (e.g. $k=d-2,d-1$ for $d=5,7,9,11$). The list of values for the coefficients $a_{k}(d)$ are summarized in Table 1. Values below $1$ have been rounded to 00. While these values are not strictly zero, the number of shots is too small to accurately assess them.
$d$ Coefficients $a_{k}(d)$
3 k=2 30 k=3 935
5 k=3 231 k=4 $10^{5}$ k=5 0
7 k=4 0 k=5 $2.3\times 10^{6}$ k=6 $1.9\times 10^{8}$ k=7 0
9 k=5 0 k=6 5 k=7 $1.1\times 10^{10}$ k=8 $3.5\times 10^{11}$ k=9 $3.7\times 10^{8}$
11 k=6 140 k=7 $2.2\times 10^{4}$ k=8 $2.6\times 10^{4}$ k=9 $3.6\times 10^{13}$ k=10 $2.1\times 10^{6}$ k=11 $5.7\times 10^{16}$
Table 1: Coefficients $a_{k}(d)$ obtained by fitting the simulation results from the realistic loss model at vanishing depolarizing noise $p_{d}=0$ (see Fig. 12a)) using Eq. (25).
When both atom loss and depolarizing noise are present (see Fig. 12b)), the logical error probability is higher and the error threshold is lower than in the simplified loss model. Nevertheless, both models exhibit qualitatively similar behavior.